# SDPose: Exploiting Diffusion Priors for Out-of-Domain and Robust Pose Estimation

### v1

@@ A.5 Details of COCO-OOD
Figure 8: COCO-OOD visualizations with Monet-style oil painting. We use CycleGAN to stylize COCO validation images into a Monet-like oil painting domain, creating an OOD split to evaluate pose estimation robustness under appearance shift.
To complement the HumanArt dataset and enable OOD evaluation under matched content and labels, we construct COCO-OOD by applying artistic style transfer to the original COCO images. We adopt the official CycleGAN framework (Zhu et al., 2017) to perform unpaired image-to-image translation from the COCO domain (natural photographs) to the target domain of Monet-style paintings. We use the monet2photo model provided in the CycleGAN repository, which is trained in an unsupervised manner to learn the mapping between Monet paintings and real-world photos. During conversion, all training and validation images in COCO are processed to produce style-transferred counterparts, while preserving their original human annotations (bounding boxes, keypoints). This yields an OOD variant of COCO in which the underlying scene structure is unchanged, but the texture, color palette, and brushstroke patterns are consistent with Monet’s artistic style. Importantly, for fair comparison and to avoid introducing priors from large-scale pretrained diffusion models, we intentionally adopt the earlier CycleGAN framework rather than more recent style transfer methods. Such stylization introduces a significant appearance shift while keeping pose-related geometric information intact, making it suitable for robust pose estimation evaluation.
@@ B.4 Full Quantitative Comparison on COCO-OOD Wholebody
Table 9: Quantitative Comparison on the COCO-OOD Wholebody validation set.
@@
Detailed whole-body pose estimation results on COCO-OOD are reported in Table 9. Breaking down by body part, SDPose matches or surpasses Sapiens on the most stable regions: body (60.0 AP vs. 59.9) and feet (62.5 vs. 63.8), and delivers a notable margin on face landmarks (+5.8 AP, 64.2 vs. 58.4), highlighting the reliability of SD-native features under appearance shifts. For hands, which are the most challenging due to fine-grained articulation and limited resolution, SDPose attains 46.3/44.9 AP on left/right hands, respectively, remaining competitive but still slightly behind Sapiens-2B (48.2/46.8). Nevertheless, the overall whole-body AP/AR of 46.6/54.8 establishes SDPose as the most robust framework under OOD whole-body evaluation.

### v2

@@ A.5 Details of COCO-OOD
Figure 9: COCO-OOD visualizations. We create an OOD split of the COCO validation set by stylizing images with CycleGAN, StyTr2, and Nano-Banana into a Monet-like oil-painting domain, enabling robustness evaluation under appearance shifts.
To complement HumanArt and enable out-of-distribution evaluation under matched content and annotations, we construct COCO-OOD (Fig. 9) by applying artistic style transfer to the original COCO images while preserving human bounding boxes and keypoints. We generate three stylistic variants, each capturing different types of artistic domain shifts. Importantly, for fair comparison and to avoid introducing priors from large-scale pretrained diffusion models, we intentionally adopt the earlier CycleGAN (Zhu et al., 2017) and StyTr2 (Deng et al., 2022) framework rather than more recent style transfer approaches.
For the Monet oil-painting variant, we employ the StyTr2 framework, which produces high-fidelity oil-painting textures and color palettes while maintaining the overall scene geometry. For the Ukiyo-e variant, we adopt the official CycleGAN implementation to translate natural photographs into the Ukiyo-e domain. Finally, we also explore a Color Sketch variant generated using the Nano-Banana model (Nano Banana, ). While Nano-Banana preserves global shape in most cases, its stylization can occasionally introduce slight pixel-level misalignment with respect to the source images. Because such spatial deviations may compromise compatibility with COCO’s keypoint annotations, we use the Color Sketch variant only as a supplemental resource for latent-space t-SNE analysis and exclude it from quantitative evaluation. By drawing from multiple style-transfer frameworks, COCO-OOD reduces the bias introduced by any single model and produces stylistic shifts with diverse characteristics. This diversity leads to a more balanced and comprehensive benchmark. Overall, COCO-OOD maintains the geometric content of COCO while producing substantial appearance shifts across distinct artistic domains, providing a controlled and annotation-preserving testbed for assessing cross-domain robustness in pose estimation.
@@ B.1 Full Quantitative Comparison on COCO
SDPose achieves 81.3 AP / 85.2 AR on the COCO validation set (Table 3) with only 40 training epochs using a 0.95B SD-v2 backbone. It matches the accuracy of Sapiens (82.1–82.2 AP) despite requiring 5× fewer epochs and a smaller backbone, and surpasses GenLoc (+3.7 AP, +4.5 AR). SDPose also outperforms ViTPose++ (+1.9 AP), which relies on multiple auxiliary datasets, while being trained solely on COCO.
@@ B.4 Full Quantitative Comparison on COCO-OOD Wholebody
Table 6: Quantitative Comparison on the COCO-OOD-Monet Wholebody validation set.
@@
Detailed whole-body pose estimation results on COCO-OOD Monet are reported in Table 6. Breaking down by body part, SDPose matches or surpasses Sapiens on the most stable regions: body (60.0 AP vs. 59.9) and feet (62.5 vs. 63.8), and delivers a notable margin on face landmarks (+5.8 AP, 64.2 vs. 58.4), highlighting the reliability of SD-native features under appearance shifts. For hands, which are the most challenging due to fine-grained articulation and limited resolution, SDPose attains 46.3/44.9 AP on left/right hands, respectively, remaining competitive but still slightly behind Sapiens-2B (48.2/46.8). Nevertheless, the overall whole-body AP/AR of 46.6/54.8 establishes SDPose as the most robust framework under OOD whole-body evaluation. Detailed whole-body pose estimation results on COCO-OOD Ukiyo-e are presented in Table 7. Breaking down by region, SDPose remains competitive with Sapiens-2B on the more structurally stable parts: body (61.2 AP vs. 62.2) and feet (64.7 vs. 67.3), and shows a clear advantage on face landmarks (+4.8 AP, 64.7 vs. 59.9), reflecting the stronger style-invariant geometry encoded by diffusion U-Net features under artistic shifts. For hands—the most challenging subset due to fine-scale pose variation—SDPose achieves 45.9/44.1 AP on left/right hands, remaining close to Sapiens-2B (49.8/46.7) despite the latter’s larger model capacity. Importantly, SDPose attains the highest whole-body AP/AR (47.7/56.4), demonstrating consistently stronger robustness than both Sapiens-1B and Sapiens-2B when evaluated under Ukiyo-e style transformations.
Table 7: Quantitative Comparison on the COCO-OOD-Ukiyoe Wholebody validation set.
Model Body AP Body AR Feet AP Feet AR Face AP Face AR
Sapiens-1B 54.0 60.9 57.9 69.9 58.7 65.3
Sapiens-2B 62.2 68.2 67.3 76.5 59.9 66.9
SDPose (Ours) 61.2 67.5 64.7 75.0 64.7 71.2
(a) Body, Feet, Face
Model L-Hand AP L-Hand AR R-Hand AP R-Hand AR Whole AP Whole AR
Sapiens-1B 43.7 52.8 40.9 50.6 40.5 49.4
Sapiens-2B 49.8 58.7 46.7 55.7 46.6 55.8
SDPose (Ours) 45.9 55.7 44.1 54.3 47.7 56.4
(b) Hands and Whole-body